I need the answers to these three questions. Please Show all work. You must write each equation in standard form. Please type answer, don't write on paper. Thanks!!!

What is the equation of a circle with a center at (-3, 6) and a radius of 7?
Answer #1

Standard form of a circle is (x-h)2 + (y-k)2 = r2 where (h,k) is center and " r " is radius of circle

1. Center (-3.6) Radius , r = 7 Equation is (x+3)2 + (y-6)2 = 72

2. Center (0,0) Radius , r = 4.5 Equation is x2 + y2 = 4.52

3. Center (1,3) Radius , r = 6 Equation is (x-1)2 + (y-3)2 = 62